# Non Null-Controllability Properties of the Grushin-Like Heat Equation on 2D-Manifolds

### v2

@@ 2 Sketch of proofs
with minimal domain $C_{c}^{\infty}(\Omega)$. We consider its Friedrich extension, still denoted by $G$, and defined by
@@ 3.2.1 Real spectral analysis
We want to estimate $(G_{V,\xi}-\mu_{\xi})v_{\xi}$ in $L^{2}$-norm, and the Proposition will follow by Lemma 3.9. Let $\varepsilon>0$ sufficiently small. First, we split the integral as
@@ 3.2.2 Complex spectral analysis
it follows that $1+V(G_{\nu}-z)^{-1}$ is an invertible linear operator on $L^{2}(I)$ for $\nu$ large enough, and so is $G_{V,\nu}-z$. Moreover,
@@ 4.1.1 Agmon estimates
which is a diffeomorphism from $(0,T)\times\Omega$ onto its image that we denote by $\mathcal{P}$ (see [darde2023null, Fig. 3.4]), and which satisfies $\mathcal{P}\subset\Omega_{x}\times U$ with (see [darde2023null, Fig. 3.4])
$U=D(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(-a)})\cup D(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(b)})\cup\{z\in\mathbb{C},|z|<1,\operatorname{arg}(z)\notin I\}.$ (116)
@@
where we recall that $\mathcal{V}$ is an arbitrary small neighborhood of $\overline{U}$ defined in (116). For every $\varepsilon>0$, for every $T>0$ such that
@@
there exists $z_{0}\in D(0,e^{-q^{\prime}(0)(1+\varepsilon)T})\setminus\overline{U}$. Hence, by Runge’s theorem, there exists a sequence of polynomials $(\tilde{p}_{k})_{k}$ that converges to $1/(z-z^{0})$, uniformly on every compact subset of $\mathbb{C}\setminus z_{0}[1,+\infty)$. We then choose the sequence of polynomials $p_{k}=z^{N+1}\tilde{p}_{k}$ to disprove (120). Indeed, as $k\rightarrow+\infty$, the left-hand side of (120) explodes, while the right-hand side remains uniformly bounded.
@@ 5.1 Reduction process
We recall that when working under H1H1-loc, our structure is of step $\gamma+1=2$, and under H1H1-glob, it is of step $\gamma+1\geq 2$. In the second part of the above Proposition, if $\gamma+1>2$, and $\varepsilon$ is such that $\omega^{\varepsilon}=\mathcal{U}$, the conclusion of our Proposition is no longer true. When the control zone $\omega^{\varepsilon}$ intersect the singularity, one needs to be careful on the regularity of the initial states in system (131) that can be steered to 0 by means of $L^{2}$-controls induced by the assumption that system (5) is null-controllable. For a structure of step $k$, one would at least be required to consider initial states in $D\left(\Delta_{\mathcal{U}}^{\frac{k-1}{2}}\right)$. However, in the proof below, some extension arguments would need to be carefully checked. Thus, in our case, and for simplicity, we restrict $\varepsilon>0$ to be small so that $\omega^{\varepsilon}$ is located within the elliptic region, and so that our argument works when $\gamma+1>2$ under H1H1-glob.
@@
$G:=S\Delta_{\mathcal{U}}S^{-1}=\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x}f)+q(x)^{2}\partial_{y}(r(y)^{2}f).$ (136)
@@
$\left\{\begin{array}[]{lcll}\partial_{t}f-(\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x})+q(x)^{2}\partial_{y}(r(y)^{2})\partial_{y})f&=&u(t,x,y)\mathbf{1}_{\omega^{\varepsilon}}(x,y),&(t,x,y)\in(0,T)\times\Omega,\\ f(t,p)&=&0,&(t,p)\in(0,T)\times\partial\Omega,\\ f(0,p)&=&f_{0}(p),&p\in\Omega,\end{array}\right$. (140)

### v3

@@ 2 Sketch of proofs
with minimal domain $C_{c}^{\infty}(\Omega)$. We consider its Friedrich extension, still denoted by $G_{V}$, and defined by
@@ 3.2.1 Real spectral analysis
We want to estimate $(G_{V,\xi}-\mu_{\xi})v_{\xi}$ in $L^{2}$-norm, and the Proposition will follow from Lemma 3.9. Let $\varepsilon>0$ sufficiently small. First, we split the integral as
@@ 3.2.2 Complex spectral analysis
it follows that $1+V(G_{\nu}-z)^{-1}$ is an invertible linear operator on $L^{2}(I)$ for $|\nu|$ large enough, and so is $G_{V,\nu}-z$. Moreover,
@@ 4.1.1 Agmon estimates
which is a diffeomorphism from $(0,T)\times\omega$ onto its image that we denote by $\mathcal{P}$, and which satisfies $\mathcal{P}\subset\Omega_{x}\times U$ with (see Figure 2)
$U$ $=D\left(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(-a)}\right)\cup D\left(0,e^{-(1-\varepsilon)d_{\operatorname{agm}}(b)}\right)\cup\{z\in\mathbb{C},|z|<1,\operatorname{arg}(z)\notin I\}$
$=D\left(0,e^{-(1-\varepsilon)\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b))}\right)\cup\{z\in\mathbb{C},|z|<1,\operatorname{arg}(z)\notin I\}.$ (116)
@@
where we recall that $\mathcal{V}$ is an arbitrary small neighborhood of $\overline{U}$ defined in (4.2). For every $\varepsilon>0$, for every $T>0$ such that
@@
there exists $z_{0}\in D(0,e^{-q^{\prime}(0)(1+\varepsilon)T})\setminus\overline{U}$ (see Figure 2). Hence, by Runge’s theorem, there exists a sequence of polynomials $(\tilde{p}_{k})_{k}$ that converges to $1/(z-z^{0})$, uniformly on every compact subset of $\mathbb{C}\setminus z_{0}[1,+\infty)$. We then choose the sequence of polynomials $p_{k}=z^{N+1}\tilde{p}_{k}$ to disprove (120). Indeed, as $k\rightarrow+\infty$, the left-hand side of (120) explodes, while the right-hand side remains uniformly bounded.
@@
Figure 2: The yellow pacman shape is $U$. The disk in red is $D\left(0,e^{-(1-\varepsilon)\min(d_{\operatorname{agm}}(-a),d_{\operatorname{agm}}(b))}\right)$, and it contains the dotted disk $D\left(0,e^{-q^{\prime}(0)(1+\varepsilon)T}\right)$. When $T$ is too small, there exists $z_{0}\in D(0,e^{-q^{\prime}(0)(1+\varepsilon)T})\setminus\overline{U}$.
@@ 5.1 Reduction process
We recall that when working under H1H1-loc, our structure is of step $\gamma+1=2$, and under H1H1-glob, it is of step $\gamma+1\geq 2$. In the second part of the above Proposition, if $\gamma+1>2$, and $\varepsilon$ is such that $\omega^{\varepsilon}=\mathcal{U}$, the conclusion of our proof of Proposition 5.3 is no longer true. Using our approach via cutoff arguments, when the control zone $\omega^{\varepsilon}$ intersects the singularity, one needs to be careful on the regularity of the initial states in system (131) that can be steered to 0 by means of $L^{2}$-controls induced by the assumption that system (5) is null-controllable. For a structure of step $k$, one would at least be required to consider initial states in $D\left(\Delta_{\mathcal{U}}^{\frac{k-1}{2}}\right)$. However, in the proof below, some extension arguments would need to be carefully checked. Thus, under H1H1-glob, for simplicity, we restrict $\varepsilon>0$ to be small so that $\omega^{\varepsilon}$ is located within the elliptic region, and so that our argument works when $\gamma+1>2$.
@@
$G:=S\Delta_{\mathcal{U}}S^{-1}=\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x})+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}).$ (136)
@@
$\left\{\begin{array}[]{lcll}\partial_{t}f-(\frac{1}{h(x)}\partial_{x}(h(x)\partial_{x})+q(x)^{2}\partial_{y}(r(y)^{2}\partial_{y}))f&=&u(t,x,y)\mathbf{1}_{\omega^{\varepsilon}}(x,y),&(t,x,y)\in(0,T)\times\Omega,\\ f(t,p)&=&0,&(t,p)\in(0,T)\times\partial\Omega,\\ f(0,p)&=&f_{0}(p),&p\in\Omega,\end{array}\right$. (140)